OpenStudy (anonymous):

A photographer conducted three sessions yesterday. She completed the first session in 56 minutes, the second in 42 minutes and the third in 1 hour and 16 minutes. How many hours did it take her to complete all three sessions?

OpenStudy (anonymous):

\[1st = 56 \min\]\[2nd = 42 \min\]\[3rd = 1 h 16 \min = 76 \min\]\[\sum = 1st + 2nd + 3rd = 56 + 42 + 76 = 174 \min = 2 h 54 \min\]

OpenStudy (anonymous):

wow thx

OpenStudy (across):

We know a photographer conducted three consecutive sessions in the following times: 56 min, 42 min, and 1 hour and 16 min. 1 hour andd 16 minutes is the same as 76 minutes. Therefore, 56 + 42 + 76 = 174 minutes = 2 hours and 54 minutes.